### 03 Mar 2026 v1

# arXiv:2603.02804v1[quant-ph] 03 Mar 2026

Abstract
While real quantum devices have been increasingly used to conduct research focused on achieving quantum advantage or quantum utility in recent years, executing deep quantum circuits or performing quantum machine learning with large-scale data on current noisy intermediate-scale quantum devices remains challenging, making classical simulation essential for quantum machine learning research. However, classical simulation often suffers from the cost of gradient calculations, requiring enormous memory or computational time. In this paper, to address these problems, we propose a method to fuse multiple consecutive gates in each of the forward and backward paths to improve throughput by minimizing global memory accesses. As a result, we achieved approximately $20$ times throughput improvement for a Hardware-Efficient Ansatz with $12$ or more qubits, reaching over $30$ times improvement on a mid-range consumer GPU with limited memory bandwidth. By combining our proposed method with gradient checkpointing, we drastically reduce memory usage, making it possible to train a large-scale quantum machine learning model, a $20$-qubit, $1,000$-layer model with $60,000$ parameters, using $1,000$ samples in approximately $20$ minutes. This implies that we can train the model on large datasets, consisting of tens of thousands of samples, such as MNIST or CIFAR-10, within a realistic time frame (e.g., $20$ hours per epoch). In this way, our proposed method drastically accelerates classical simulation of quantum machine learning, making a significant contribution to quantum machine learning research and variational quantum algorithms, such as verifying algorithms on large datasets or investigating learning theories of deep quantum circuits like barren plateau.
keywords:
[…]
1 Introduction
In recent years, quantum machine learning (QML), such as quantum neural networks [farhi2018classification, mitarai2018quantum], and variational quantum algorithms (VQAs) [cerezo2021variational], such as variational quantum eigen-solver (VQE) [peruzzo2014variational, kandala2017hardware] and quantum approximate optimization algorithm (QAOA) [farhi2014quantum], have been actively researched. Current noisy intermediate-scale quantum devices have limitations in the depth of quantum circuits and the connectivity of qubits, so classical simulation is still important to verify new algorithms or investigate learning theories, including barren plateaus [mcclean2018barren, wang2021noise, cerezo2021cost] and the existence of many local minima [you2021exponentially, bittel2021training], particularly in deep quantum circuits.
Most of the current research on classical simulation of quantum circuits [pednault2019leveraging, tindall2024efficient] focuses on simulating a single large quantum circuit, supporting the achievement of quantum advantage [arute2019quantum, google2025observation] or quantum utility [kim2023evidence]. In general, a classical simulation using a state vector [suzuki2021qulacs, qiskit2024] requires exponentially large memory with respect to the number of qubits, so it is usual to utilize a supercomputer or multi-GPU clusters [pednault2019leveraging, de2019massively, kim2025scaleqsim, haner20175, smelyanskiy2016qhipster], or to use storage devices [pednault2019leveraging, park2022snuqs] for classical simulation over $30$ qubits. The existing classical simulators are superior at computing a single large quantum circuit, but classical simulation of QML requires memory that grows linearly with batch size, as well as exponentially with the number of qubits, making it challenging even for a $20$-qubit quantum circuit. Moreover, there is a challenge in calculating gradients [mitarai2018quantum, jones2020efficient], as it requires enormous computation time or memory. However, most of the existing general-purpose quantum circuit simulators are limited to optimizing a single quantum circuit, or are not specifically designed for batch processing, often failing to take advantage of the parallel performance of GPUs due to inefficient memory access.
A popular approach for gradient calculation is the parameter-shift rule [mitarai2018quantum]. This approach is memory-efficient, but the number of quantum circuit executions scales linearly with the number of variational gates, making it impractical for large-scale models in classical simulation. On the other hand, the adjoint method is fast because the effective number of quantum circuit executions is independent of the number of variational gates. However, a practical implementation of the adjoint method faces a critical trade-off between memory usage and execution time regarding the management of intermediate states. A naive implementation stores all intermediate states of quantum gates in the forward path in order to reuse them in the backward path, causing memory usage to increase linearly with the number of gates. Alternatively, a memory-efficient implementation avoids storing intermediate states by sequentially applying each gate to recompute them during the backward path [jones2020efficient, luo2020yao], leading to a significant increase in computational time due to the repeated accesses to exponentially large quantum states [jones2020efficient, luo2020yao]. Therefore, existing simulators typically optimize only forward paths using gate fusion [suzuki2021qulacs, haner20175, smelyanskiy2016qhipster], or sequentially apply quantum gates to obtain intermediate states in the backward paths [jones2020efficient, luo2020yao], leaving the backward path bottlenecked by inefficient memory accesses.
In this paper, we propose a method to fuse multiple consecutive quantum gates into a single fused operator for each of the forward and backward paths, without writing any intermediate states of the fused gates to global memory. This approach allows us to reduce the memory requirements for storing quantum states and decrease the number of accesses to global memory, leading to an enhancement of the arithmetic intensity. However, if we simply read the input state and temporarily store all recomputed intermediate states, the number of variables to be held in registers can easily exceed the register capacity. This causes register spilling and increases the computation time. Therefore, to reduce this register pressure, we minimize the number of intermediate states held in registers, and load unitary matrices whenever necessary. As a result, our proposed method achieves approximately $20$ times improvement in throughput compared to a PyTorch native implementation for a typical single-layer Hardware-Efficient Ansatz (HEA) [kandala2017hardware]. In particular, for a mid-range GPU with limited memory bandwidth, we achieved over $30$ times improvement in throughput, highlighting the effectiveness of our proposed method on GPUs with limited memory bandwidth.
Moreover, since our implementation uses Triton [triton2019] and PyTorch [Ansel_PyTorch_2_Faster_2024] and is integrated with PyTorch’s automatic differentiation, it is easy to utilize the existing PyTorch ecosystem. For example, by combining our proposed method with gradient checkpointing, we can reduce memory usage to $O(\sqrt{d})$ for $d$ HEA layers. In our numerical experiment, we performed a task of minimizing a sum of batched expectation values for an observable at the final states of a $20$-qubit with $1,000$ HEA layers ($60,000$ parameters). As a result, we succeeded in running $1,000$ samples in approximately $20$ minutes per epoch. This result indicates that we can train a model consisting of $60,000$ parameters using tens of thousands of samples within $20$ hours per epoch, i.e., a realistic time frame. Further acceleration of training can be easily realized using multiple GPUs with PyTorch’s distributed training utilities, such as DistributedDataParallel.
In this way, since our proposed method enables the training of QML with large datasets or deep quantum circuits within a realistic timeframe, it is expected to significantly accelerate research and development in QML.
2 Preliminaries
In this section, we explain the adjoint method for efficiently calculating gradients.
2.1 State vector simulation and objective function
Here, we consider a quantum state of an $n$-qubit system represented by a complex vector of length $2^{n}$, $|\psi\rangle\in\mathbb{C}^{2^{n}}$, and a parameterized quantum circuit (PQC) consisting of a series of $M$ parameterized unitary gates $U(\bm{\theta})=U_{M}(\theta_{M})\cdots U_{1}(\theta_{1})$ characterized by trainable parameters $\bm{\theta}=\{\theta_{1},\dots,\theta_{M}\}$. We represent the initial state as $|\psi_{0}\rangle$, the intermediate state as $|\psi_{j}\rangle\coloneqq U_{j}\ldots U_{1}|\psi_{0}\rangle$, and the final state as $|\psi_{M}(\bm{\theta})\rangle\coloneqq U(\bm{\theta})|\psi_{0}\rangle$.
As is often the case in QML and VQAs, we focus on the task of minimizing a loss function $L$ by optimizing the trainable parameters $\bm{\theta}$. Let $O$ denote an observable. We represent the expectation value $E(\bm{\theta})$ as follows:
$E(\bm{\theta})=\langle\psi_{M}(\bm{\theta})|O|\psi_{M}(\bm{\theta})\rangle.$ (1)
Although the loss function $L$ is a function of $E(\bm{\theta})$ in general, we assume $L=E$ in our benchmarks for simplicity.
2.2 Adjoint method
As we mentioned in Sec. 1, the adjoint method can efficiently calculate gradients in classical simulation. In this subsection, we explain the adjoint method for optimizing trainable parameters $\bm{\theta}$ in a PQC, $U(\bm{\theta})$. Specifically, we explain how to calculate gradients for the $j$-th unitary gate $U_{j}(\theta_{j})$.
In the adjoint method, we introduce an adjoint state $|\lambda_{j}\rangle$ for the backward path. For the $j$-th unitary gate $U_{j}(\theta_{j})$, the adjoint state is propagated by
$|\lambda_{j-1}\rangle=U_{j}^{\dagger}(\theta_{j})|\lambda_{j}\rangle.$ (2)
Using this adjoint state $|\lambda_{j}\rangle$, the gradient of the loss function $L$ with respect to $\theta_{j}$ can be described by
$\frac{\partial L}{\partial\theta_{j}}=\text{Re}\left\langle\lambda_{j}\left|\frac{\partial U_{j}}{\partial\theta_{j}}\right|\psi_{j-1}\right\rangle.$ (3)
Note that from Eq. (3), we need both the input quantum state $|\psi_{j-1}\rangle$ and the adjoint state $|\lambda_{j}\rangle$ to calculate the gradient with respect to each trainable parameter $\theta_{j}$. This indicates that we need to preserve the input quantum states $|\psi_{j-1}\rangle$ to efficiently compute the gradients. Namely, a naive implementation requires $O(M\cdot 2^{n})$ memory to store all the input quantum states $\{|\psi_{0}\rangle,\dots,|\psi_{M-1}\rangle\}$ in memory. As the number of variational parameters $M$ grows, the required memory can easily exceed the memory capacity, making classical simulation infeasible. Even within the memory capacity, the computation time significantly increases due to repeated accesses to these exponentially large vectors in global memory. We will tackle these problems with our proposed method in the following sections.
Also note that, since we assume $L=E$, the adjoint state at the final gate can be described by
$|\lambda_{M}\rangle=2O|\psi_{M}\rangle.$ (4)
3 Methods
In this section, we describe the implementation of classical simulation of a quantum circuit, focusing specifically on the efficient usage of GPU memory. Since the performance of classically simulating a quantum circuit is typically limited by the memory bandwidth, we explain a kernel fusion technique to reduce the data transfer, and a recomputation technique to prevent register spilling.
3.1 Applying a quantum gate
As we mentioned in Sec. 2.1, an $n$-qubit state vector $|\psi\rangle$ is represented by a complex vector of length $2^{n}$. We describe $|\psi\rangle$ using computational basis states $|k\rangle$ and complex probability amplitudes $a_{k}$, whose index $k$ $(0\leq k<2^{n})$ corresponds to its binary representation $b_{n-1}\ldots b_{0}$ $(k=\sum_{i=0}^{n-1}b_{i}2^{i})$ of length $n$, as follows:
$|\psi\rangle=\sum_{k=0}^{2^{n}-1}a_{k}|k\rangle=a_{0\ldots 00}|0\cdots 00\rangle+a_{0\ldots 01}|0\cdots 01\rangle+\cdots+a_{1\ldots 10}|1\cdots 10\rangle+a_{1\ldots 11}|1\cdots 11\rangle$
Next, we consider applying a single qubit gate $u$ on the $t$-th qubit, where $u$ is represented by
$u=\begin{pmatrix}u_{00}&u_{01}\\ u_{10}&u_{11}\end{pmatrix}$,
and $U$ is represented by $U=I\otimes\cdots\otimes\underbrace{u}_{t\text{-th}}\otimes\cdots\otimes I$.
This operation can be described by a linear transformation for pairs of probability amplitudes ($a_{\ldots 0_{t}\ldots}$, $a_{\ldots 1_{t}\ldots}$), whose indexes only differ at the $t$-th bit, updating them to ($a^{\prime}_{\ldots 0_{t}\ldots}$, $a^{\prime}_{\ldots 1_{t}\ldots}$) as follows:
$\begin{pmatrix}a^{\prime}_{...0_{t}...}\\ a^{\prime}_{...1_{t}...}\end{pmatrix}=\begin{pmatrix}u_{00}&u_{01}\\ u_{10}&u_{11}\end{pmatrix}\begin{pmatrix}a_{...0_{t}...}\\ a_{...1_{t}...}\end{pmatrix}$,
where the bits in the dotted parts of the binary representation remain unchanged. In a parallelized implementation on a GPU, each thread typically loads a pair of amplitudes from global memory, applies the unitary matrix, and writes the resulting amplitudes back to global memory. This operation is performed for all of the $2^{n-1}$ pairs, requiring a huge number of global memory accesses.
Next, we explain how to update amplitudes when applying two-qubit gates. First, when applying a CZ gate acting on control qubit $c$ and target qubit $t$ to a quantum state, we only need to multiply the amplitudes $a_{\ldots 1_{c}\ldots 1_{t}\ldots}$ (or $a_{\ldots 1_{t}\ldots 1_{c}\ldots}$) by 1-1, where the control and target bits of the index are both $1$.
Similarly, when applying CNOT gate acting on control qubit $c$ and target qubit $t$ to a quantum state, we can update the quantum state by swapping the pair of amplitudes $a_{\ldots 1_{c}\ldots 0_{t}\ldots}$ and $a_{\ldots 1_{c}\ldots 1_{t}\ldots}$ (or $a_{\ldots 0_{t}\ldots 1_{c}\ldots}$ and $a_{\ldots 1_{t}\ldots 1_{c}\ldots}$), i.e., swapping the amplitudes corresponding to the $t$-th bit being 0 and $1$ wherever the $c$-th bit is $1$.
Therefore, whenever we apply a quantum gate, we must read and write $O(2^{n})$ probability amplitudes, which limits the performance of classical simulation on GPUs due to memory bandwidth constraints.
3.2 Gate fusion in the forward path
To save memory bandwidth in the forward path, a popular technique is to fuse multiple quantum gates before applying them to a quantum state [suzuki2021qulacs, haner20175, smelyanskiy2016qhipster]. In this paper, to maintain consistency with the calculation in the backward path, we consider applying $m$ consecutive single-qubit gates, each denoted as $u^{(1)},\ldots,u^{(m)}$, acting on the same target qubit $t$, such that:
$\begin{pmatrix}a^{\prime}_{...0_{t}...}\\ a^{\prime}_{...1_{t}...}\end{pmatrix}$ $=u^{(m)}\cdots u^{(1)}\begin{pmatrix}a_{...0_{t}...}\\ a_{...1_{t}...}\end{pmatrix}$
$=\underbrace{\begin{pmatrix}u_{00}^{(m)}&u_{01}^{(m)}\\ u_{10}^{(m)}&u_{11}^{(m)}\end{pmatrix}\cdots\begin{pmatrix}u_{00}^{(1)}&u_{01}^{(1)}\\ u_{10}^{(1)}&u_{11}^{(1)}\end{pmatrix}}_{\text{Fused Unitary}}\begin{pmatrix}a_{...0_{t}...}\\ a_{...1_{t}...}\end{pmatrix}$.
By performing this operation, we can update a quantum state by reading and writing its state vector only once per fused gate consisting of $m$ consecutive single-qubit gates. This operation allows us to reduce memory accesses by a factor of $m$. To accelerate the classical simulation of typical Ansatzes in QML and VQA, in this paper, we consider applying a fused gate consisting of consecutive single-qubit gates acting on the same qubit and also the adjacent qubits.
Since the adjoint method requires the input state vector for each variational gate, our implementation stores only the output state vector of each fused variational gate during the forward path. As detailed in the next subsection, we then use this stored output state vector to recompute the input state vector for each constituent gate within the fused operator.
When storing these state vectors, we can store their real and imaginary parts with original precision (float32 or float64), or with reduced precision (bfloat16 or float32) to further reduce the memory usage and use larger batch sizes. We refer to this reduced-precision approach as the memory-saving mode (denoted as Triton fused (mem save) in our numerical experiments). Even when we store them with halved precision, we perform the calculations in original high precision using (float32 or float64), i.e., the precision is only reduced when storing them in global memory. This technique allows us to mitigate the impact on computational accuracy while halving memory usage. Furthermore, in our numerical experiments, we will demonstrate the efficient training of deep quantum circuits by reducing the storage of state vectors using our proposed method combined with the gradient checkpointing technique.
3.3 Gate fusion in the backward path
Next, we explain the gate fusion technique for the backward path. One of the major contributions of our work is the realization of efficient gate fusion in the backward path.
As shown in Eq. (3), when we calculate the gradients of variational parameters $\theta_{j+k}$, we need both the input quantum state $|\psi_{j+k-1}\rangle$ and the adjoint state $|\lambda_{j+k}\rangle$ for the corresponding variational gate. If we simply fuse $m$ consecutive variational gates in the backward path, it requires storing the elements of quantum states $|\psi_{j}\rangle,\ldots,|\psi_{j+m-1}\rangle$, $|\lambda_{j}\rangle,\ldots,|\lambda_{j+m}\rangle$ in registers. Therefore, as $m$ increases, the required memory can easily exceed the register capacity of the GPU, causing them to be spilled to a low-speed memory region, which increases computation time.
Therefore, by focusing on the reversibility of quantum gates, we recompute the intermediate states when needed, instead of storing all of them. Specifically, as shown in Fig. 1, given the quantum state $|\psi_{j+m}\rangle$ (stored during the forward path) and the adjoint state $|\lambda_{j+m}\rangle$ as inputs for the fused gates consisting of $m$ consecutive variational gates, we perform the following procedure to calculate the gradients.
  • Recomputing the input quantum states for $U_{j+k}(\theta_{j+k})$ and calculating the gradients $\frac{\partial L}{\partial\theta_{j+k}}$ for $k=1,\ldots,m$
    We calculate the gradients, input quantum states, and adjoint states for $U_{j+k}(\theta_{j+k})$ by iterating the gate index of the fused gates in reverse order from $k=m$ to $k=1$. For each step $k$,
    1. 1.
      Updating quantum states:
      We load $U_{j+k}(\theta_{j+k})$ and apply it to $|\psi_{j+k}\rangle$ in order to calculate $|\psi_{j+k-1}\rangle$:
      $|\psi_{j+k-1}\rangle=U_{j+k}^{\dagger}(\theta_{j+k})|\psi_{j+k}\rangle$
    2. 2.
      Calculating gradients:
      For a variational single-qubit rotation gate, we can easily calculate $\frac{\partial u_{j+k}(\theta_{j+k})}{\partial\theta_{j+k}}=-\frac{1}{2}\sin{\frac{\theta_{j+k}}{2}}I-i\frac{1}{2}\cos{\frac{\theta_{j+k}}{2}}P$ from the definition $u_{j+k}(\theta_{j+k})=\cos{\frac{\theta_{j+k}}{2}}I-i\sin{\frac{\theta_{j+k}}{2}}P$, where $P\in\{I,X,Y,Z\}$.
      According to Eq. (3), we calculate the gradient $\frac{\partial L}{\partial\theta_{j+k}}$.
    3. 3.
      Updating adjoint state:
      We update the adjoint state using the loaded gate $U_{j+k}(\theta_{j+k})$:
      $|\lambda_{j+k-1}\rangle=U_{j+k}^{\dagger}(\theta_{j+k})|\lambda_{j+k}\rangle$
By repeating this procedure, we can efficiently calculate all the gradients of the training parameters $\bm{\theta}$.
Using this procedure, each thread only needs to store a part of the elements of two exponentially large vectors ($|\psi_{j+k-1}\rangle$ and $|\lambda_{j+k}\rangle$) in registers during kernel processing. To avoid register spilling, we read $U_{j+k}(\theta_{j+k})$ when required. Compared to the method of storing all the input quantum states and adjoint states, although our method requires additional calculations to recompute the input quantum states for each $U_{j+k}(\theta_{j+k})$, it allows us to mitigate the bottleneck, i.e., decrease the accesses to global memory. This is expected to reduce the required memory and improve throughput by reducing the number of stored quantum states.
Figure 1: Conventional method and proposed method
3.4 Theoretical memory analysis with gradient checkpointing
By combining our proposed method with gradient checkpointing (specifically using PyTorch’s torch.utils.checkpoint), we can significantly reduce the memory usage, especially for deep quantum circuits.
To theoretically analyze the behavior of memory usage, we consider a typical deep parameterized quantum circuit consisting of a repeating layer structure, such as a Hardware-Efficient Ansatz (HEA). Let $d$ be the total number of HEA layers in a quantum circuit, and let $l_{\text{var}}$ and $l_{\text{const}}$ represent the number of variational gates and constant gates per layer, respectively.
In the naive implementation of the adjoint method without gradient checkpointing, since we need to store the input quantum states of all gates for the backward path, the number of required state vectors to be stored is $(l_{\text{var}}+l_{\text{const}})d$. To mitigate this linear memory growth, we apply gradient checkpointing by dividing the quantum circuit into checkpoint blocks, each consisting of $b$ consecutive HEA layers, as illustrated in Fig. 2. For simplicity, we assume that $d$ is divisible by $b$.
When combining the naive implementation with gradient checkpointing, the number of required state vectors to be stored for the backward path becomes $(l_{\text{var}}+l_{\text{const}})b+d/b$. The first term represents the stored intermediate states within a single active checkpoint block, and the second term ($d/b$) arises from the stored input quantum states for each checkpoint block, which are required to recalculate the intermediate state vectors within the checkpoint block during gradient checkpointing. Therefore, letting $M_{\text{sv}}$ be the required memory for a single state vector, the required memory $M_{total}$ (bytes) for the PyTorch native implementation is estimated as
$M_{total}(b)$ $\simeq\left((l_{\text{var}}+l_{\text{const}})\cdot b+\frac{d}{b}\right)\times M_{\text{sv}}$ (5)
$\geq 2\sqrt{(l_{\text{var}}+l_{\text{const}})d}\times M_{\text{sv}},$ (6)
where the equality holds when $b^{*}=\sqrt{\frac{d}{l_{\text{var}}+l_{\text{const}}}}$.
Similarly, for our proposed method combined with gradient checkpointing, we must store $d/b$ input quantum states for the checkpoint blocks. Additionally, within the active checkpoint block, we only need to store $\left\lceil\frac{l_{\text{var}}}{m}\right\rceil\cdot b$ output state vectors of the fused variational gates, since the implementation of our proposed method does not require storing any input and output state vectors for constant gates. By introducing $\alpha$ as an effective number of fused gates to uniformly treat the memory reduction effects regardless of the type of saved tensors, the required memory (bytes) can be estimated as
$M_{total}(b)$ $\simeq\left(\left\lceil\frac{l_{\text{var}}}{\alpha}\right\rceil\cdot b+\frac{d}{b}\right)\times M_{\text{sv}}$ (7)
$\geq 2\sqrt{\left\lceil\frac{l_{\text{var}}}{\alpha}\right\rceil d}\times M_{\text{sv}},$ (8)
where the equality holds when $b^{*}=\sqrt{d/\left\lceil\frac{l_{\text{var}}}{\alpha}\right\rceil}$. Specifically, for our proposed method (Triton fused), $\alpha$ corresponds to $m$. For our proposed method with memory-saving mode (Triton fused (mem save)), we can effectively regard $\alpha=2m$ because we store the real and imaginary parts of the state vector as bfloat16 for complex64 or float32 for complex128 to halve the memory required to store state vectors. Note that this memory-saving technique applies only to the output state vectors stored for each fused variational gate within an active checkpoint block. The input quantum states for checkpoint blocks are stored without type conversion because they are automatically managed by PyTorch’s checkpointing mechanism.
Consequently, by combining our proposed method with gradient checkpointing, the memory usage scales as $O(\sqrt{d})$. This combined approach allows us to drastically reduce the total memory requirements while maintaining the high throughput achieved by minimizing global memory accesses, enabling efficient classical simulation of deep quantum circuits.
Figure 2: Stored state vectors when combining our proposed method with gradient checkpointing
4 Numerical experiments
In this section, to evaluate the performance of our proposed method, implemented with Triton [triton2019] and PyTorch [Ansel_PyTorch_2_Faster_2024], we measured the average execution time over $1,000$ runs using the benchmarking utility triton.testing.do_bench in Triton [triton2019]. We also measured peak memory usage using torch.cuda.max_memory_allocated from a single measurement, since peak memory usage remains consistent across multiple executions. Note that although our implementation supports both single precision (float32/complex64) and double precision (float64/complex128), we use float32 (complex64) in our numerical experiments to focus on performance evaluation for practical QML tasks, which often have severe memory constraints. Furthermore, throughout our numerical experiments, we apply quantum circuits to randomly initialized quantum states $|\psi_{0}\rangle$ rather than $|0\rangle^{\otimes n}$ in order to avoid artificial speedups caused by sparsity and to reflect the actual computational load and memory bandwidth utilization.
4.1 The effect of gate fusion
In this subsection, we investigate the effect of the gate fusion on a quantum circuit consisting of consecutive single-qubit gates, as shown in Fig. 3. Using a $20$-qubit system and a batch size of $100$, we focus on how memory usage and execution time depend on the number of fused gates in the forward and backward paths.
First, we measured the mean execution time per fused gate for a quantum circuit where $m$ consecutive single-qubit gates (Rx, Ry, and Rz) are applied to each qubit during the forward path. As shown in Fig. 4(a), we can see that our proposed method allows us to drastically accelerate the classical simulation, compared with the PyTorch native implementation, which applies each gate sequentially. In our numerical experiments, since we apply the gates in the order of Rx, Ry, and Rz, the execution time of the PyTorch native implementation gradually increases, especially when applying the diagonal matrix Rz gate. Furthermore, Fig. 4(a) shows that the execution time of our proposed method increases almost linearly with $m$, but much more gradually than that of the PyTorch native implementation.
As shown in Fig. 4(b), the peak memory usage is almost the same regardless of the number of gates to be fused. In the forward path, the peak memory usage of our proposed method (Triton fused) is the lowest among the three implementations. The total peak memory usage is $2,500$ MB, roughly corresponding to three batched state vectors ($3\times 800$ MB). It appears that the input, output, and temporary batched state vectors are held, and the remaining $100$ MB is internal overhead. Similarly, the Triton fused (mem_save) implementation used an additional $400$ MB, caused by storing the batched state vector converted to bfloat16 for the backward path. The PyTorch native implementation consumes approximately $800$ MB more memory than the Triton fused implementation, corresponding to a single batched state vector. This suggests that the PyTorch native implementation temporarily keeps a copy of the batched state vectors during the forward path, likely due to explicitly cloning them for safety. Since PyTorch’s automatic differentiation requires out-of-place operations on inputs to preserve the computation graph, our implementation is considered to be memory-efficient.
Figure 3: A quantum circuit consisting of $m$ consecutive Rx, Ry, and Rz gates used for the benchmark of verifying the effect of our proposed method
(a) Execution time per fused gate
(b) Total peak memory usage
Figure 4: (a) Execution time per fused gate and (b) Total peak memory usage when applying $m$ consecutive single-qubit gates on each qubit during the forward path
Next, to evaluate the performance of the backward function itself, we apply the same quantum circuit as above, shown in Fig. 3, and measure the mean execution time per fused kernel. From the results shown in Fig. 5(a), for consumer GPUs such as the RTX 5070 and RTX 4090, the computation time of applying a few gates to the input quantum state is almost constant, i.e., the recomputation time is negligible because the memory access latency to read quantum states and adjoint states from global memory is dominant. As the number of gates to be fused further increases, the computation time begins to scale linearly. On the other hand, for data center GPUs with large memory bandwidth such as the GH200, the computation time increases linearly with the number of gates to be fused from the beginning. As shown in Fig. 5(b), the peak memory usage linearly increases with the number of gates to be fused, and only $30$ KB is increased (from $2516.585$ MB to $2516.613$ MB). This indicates that the computation of the fused gates in the backward path is processed efficiently while mitigating the use of temporary variables.
(a) Execution time per fused gate
(b) Total peak memory usage
Figure 5: (a) Execution time per fused gate and (b) Total peak memory usage when applying $m$ consecutive single-qubit gates on each qubit at the backward path
Finally, we measured the mean execution time per fused gate and total peak memory usage over the forward and backward paths while performing a task of minimizing a sum of expectation values for the observable $O=IXYZ\ldots IXYZ$ (repeated $5$ times) using the same quantum circuit as above, shown in Fig. 3, with $20$ qubits and a batch size of $100$. The mean execution time per fused gate is shown in Fig. 6(a) and the total peak memory usage over $20$ fused gates is shown in Fig. 6(b). We performed the evaluation of the PyTorch native implementation for this numerical experiment using only GH200 because it requires a huge amount of memory to store the state vectors. Also, we evaluated our proposed method on the RTX 4090 and GH200 due to the insufficient memory capacity on the RTX 5070.
As shown in Fig. 6(a), the implementation of our proposed method performs significantly faster than the PyTorch native implementation. The mean execution time per fused gate for both approaches increases linearly with the number of variational gates to be fused, but that for our proposed method increases more gradually. This trend suggests that reducing the number of accesses to global memory by our proposed method contributes to this acceleration. Additionally, for the result of PyTorch native implementation, the increase in execution time is gradual when an Rz gate is added. As shown in Fig. 6(b), the peak memory usage of the PyTorch native implementation shows a steep linear increase with the number of variational gates to be fused, since all input quantum states for the gates are stored during the forward path for later use in the backward path. For our proposed method, on the other hand, the required memory remains almost constant regardless of the number of variational gates to be fused, and is drastically lower than that of the PyTorch native implementation, since it only needs to store the resulting state of the fused variational gates. In addition, Triton fused (mem save) implementation further reduced the peak memory usage by type converting the state vector from float32 to bfloat16 when temporarily storing it in the forward path. As shown in Fig. 6(c), when $10$ or more variational gates are fused, the mean execution time per fused gate of Triton fused (mem save) is $5\%$ to $10\%$ slower than that of Triton fused on the GH200, and $10\%$ to $20\%$ slower on the RTX 4090. Importantly, we can reduce peak memory usage by $40\%$ at the cost of an approximate $10\%$ increase in computation time. This trade-off is practically important, particularly for a consumer GPU with limited memory capacity and bandwidth.
(a) Execution time per fused gate
(b) Total peak memory usage
(c) Execution time ratio
Figure 6: (a) Execution time per fused gate and (b) total peak memory usage (c) execution time ratio between Triton Fused and Triton Fused (mem save) when minimizing the sum of expectation values at the final quantum state after applying the fused single-qubit gates shown in Fig. 3 to a $100$ fixed batch of random initial states for one epoch.
4.2 Throughput for a Hardware Efficient Ansatz
Here, we measure the throughput for a task of minimizing a sum of expectation values for the observable $O=IXYZ\cdots IXYZ$ (repeating $IXYZ$ until the length matches the number of qubits) at the final state, made by applying an HEA (shown in Fig. 7 with $d=1$ ), consisting of Rx, Ry, Rz, and CZ gates, to random initial quantum states. To evaluate the throughput, we set the batch size as large as possible to be stored within the GPU memory, and we estimated the throughput using this batch size and the mean execution time across $1,000$ runs. The results are shown in Fig. 8.
In this numerical experiment, to ensure a sufficient number of gates to be fused, we fused the consecutive single-qubit gates acting on not only the same qubit but also on adjacent qubits, as shown in Fig. 7. This is because fusing gates acting only on the same qubit limits the number of gates that can be fused, potentially making it difficult to hide memory access latency and thereby limiting the acceleration effect. Similarly, we also fused consecutive CZ gates.
As shown in Fig. 8 and Fig. 10, for $12$ qubits and above, our proposed method achieved approximately $20$ times higher throughput compared to the PyTorch native implementation. In particular, on the midrange GPU RTX5070 with limited memory bandwidth, our proposed method achieved approximately $30$ times higher throughput. These results demonstrate that implementing gate fusion in both forward and backward paths yields a significantly higher throughput than the PyTorch native implementation.
For reference, we also plot the results of an existing QML library TorchQuantum [torchquantum2024], with a red dotted line in Fig. 8. The implementation of this library uses a matrix-based approach (e.g., for calculating expectation values), making classical simulation for over $15$ qubits infeasible due to memory constraints. On the other hand, the implementation of our proposed method employs a matrix-free approach, similar to the fast quantum circuit simulator “Qulacs" [suzuki2021qulacs]. Furthermore, by recomputing the intermediate states of fused gates during the backward path, we do not need to store them in global memory. As a result, we achieved a throughput of $3.57$ samples/sec for a $29$-qubit HEA by Triton fused on a single GPU GH200 and $1.62$ samples/sec for a $30$-qubit HEA by Triton fused (mem save) on the same GPU, whereas the Triton fused implementation failed at this scale due to memory constraints. This highlights the practical effectiveness of our memory-saving strategy for large-scale classical simulations.
Figure 7: HEA and gate fusion method: the quantum circuit consists of single-qubit gates (Rx, Ry, Rz) and CZ gates. The dotted frames indicate the range of the gate fusion in our proposed method. We fused up to nine single-qubit gates, including single-qubit gates acting on the same and adjacent qubits, and also the consecutive CZ gates in the implementation of our proposed method. In our numerical experiments, we used a quantum circuit consisting of $d$ layers of this HEA.
Figure 8: Throughputs for $d=1$ of the HEA shown in Fig. 7
4.3 Classical simulation of $1,000$ layers of HEA using gradient checkpointing
Finally, as a practical benchmark, we perform a large-scale classical simulation of a deep quantum circuit for QML by combining our proposed method with gradient checkpointing in PyTorch (torch.utils.checkpoint) on the GH200. Specifically, we perform the classical simulation of minimizing the sum of expectation values for the observable $O=IXYZ\cdots IXYZ$ ( consisting of $5$ repetitions of $IXYZ$ ) at the final states, generated by applying a $20$-qubit HEA with $d=1,000$ layers (including $60,000$ parameters), as shown in Fig. 7, to random initial quantum states.
4.3.1 Preliminary experiment to determine checkpoint block size $b$
First, to optimize the checkpoint block size $b$, we measured the average execution time and peak memory usage with a batch size of one. The results are shown in Fig. 9.
Analysis of peak memory usage: As shown in Fig. 9(a), for a small $b$ ($1\leq b\leq 5$), all implementations consume a large amount of memory. This is because a small block size implies a large total number of blocks ($d/b$), making the storage of input state vectors for each block dominant, scaling as $O(d/b)$. As $b$ increases, peak memory usage begins to decrease. Then, the minimum peak memory usage for Triton fused is $1,443$ MB at $b=10$, and that for Triton fused (mem save) is $1,032$ MB at $b=20$. This result demonstrates that we successfully reduced peak memory usage by approximately $30\%$ by temporarily storing the output state vectors of fused variational gates as bfloat16. As $b$ increases further, the memory required to store output state vectors for fused gates within each active block becomes dominant, causing the peak memory usage to increase again, scaling as $O(\lceil\frac{l_{\text{var}}}{\alpha}\rceil\cdot b)$. Interestingly, at $b=100$, the peak memory usage for Triton fused is $5,972$ MB, and that for Triton fused (mem save) is $3,045$ MB, which is almost half the size. This result indicates that when the term $\left(\left\lceil\frac{l_{\text{var}}}{\alpha}\right\rceil\cdot b\right)$ dominates, the effect of reducing peak memory usage by storing compressed state vectors is consistent with theoretical expectations.
Comparison with theoretical analysis: Here, we compare the peak memory usage of our numerical results with the theoretical analysis. To determine $\alpha$, we first focus on the number of fused gates and stored state vectors. As shown in Fig. 7, for three adjacent qubits, we fused the sequence of Rx, Ry, and Rz gates acting on each of the three adjacent qubits into a single operator ($m=3\times 3=9$). For the remaining two qubits, we fused the corresponding sequence of gates ($m=2\times 3=6$). So, the number of fused variational gates is $\lceil\frac{l_{\text{var}}}{\alpha}\rceil=\lceil\frac{60}{9}\rceil=7$, and each fused variational gate requires storing its output state vector. Note that a fused CZ gate does not require storing its state vectors because it does not have a variational parameter. Since the size of complex64 is $8$ bytes, the theoretical peak memory usage for a $1,000$-layer HEA can be estimated by substituting $n=20$ and $d=1000$ into Eq. (7):
  1. 1.
    For the Triton fused implementation, it is given by
    $M_{total}=(7b+1000/b)\times 8\times 2^{20}.$ (9)
  2. 2.
    For the Triton fused (mem save) implementation, the coefficient for storing the output state vectors of fused variational gates in each active block is effectively halved because they are stored as bfloat16 from float32,
    $M_{total}=(3.5b+1000/b)\times 8\times 2^{20}.$
From Fig. 9(a), we can see that the measured values agree well with the theoretical values. In addition, the optimal block size to minimize peak memory usage is $b^{*}=\sqrt{\frac{1000}{7}}\simeq 12$ for the Triton fused implementation, and $b^{*}=\sqrt{\frac{1000}{3.5}}\simeq 17$ for Triton fused (mem save) implementation. These values are close to the measured values ($10$ and $20$, respectively). This consistency means that the peak memory usage behavior of our proposed method is highly predictable, facilitating the estimation of required resources, even for large-scale models.
Execution time and scalability: From Fig. 9(b), the execution time remains almost constant regardless of $b$ within a $1\%$ fluctuation. This result indicates that the cost of recomputation is almost hidden by the memory access latency. These slight fluctuations in the execution time are caused by the overhead of calling the gradient checkpointing function, as well as the recomputation, storage, and type conversion of state vectors within each active block during gradient checkpointing. When $b$ is small (i.e., many checkpoints), it is considered that the overhead of calling the gradient checkpointing function causes a slight increase in execution time. When $b$ is large, the computational cost of recomputing and storing state vectors in each active block grows, causing an increase in execution time. Moreover, for Triton fused (mem save), the type conversion further increases execution time. Notably, the optimal block size $b$ for the fastest execution time for the Triton fused (mem save) implementation is smaller than that for Triton fused. This is because the input state vectors for each checkpoint block during gradient checkpointing are stored without type conversion, whereas the output state vectors of fused variational gates within an active checkpoint block require type conversion, making the optimal block size $b$ for Triton fused (mem save) smaller in terms of total execution time.
4.3.2 Numerical experiment for a practical scale QML benchmark
Finally, assuming the training of a practical QML model, we measured the mean training time per epoch and peak memory usage for $1,000$ samples of randomly initialized states. For the numerical settings, we set $b=10$ for Triton fused and $b=20$ for Triton fused (mem save), based on the previous numerical results with a batch size of one. In addition, we set the mini-batch size to $60$ for Triton fused and $85$ for Triton fused (mem save), to fully use the GPU memory capacity, reaching a peak memory usage of approximately $95$ GB, as shown in Table 1.
As a result, the execution time was $1,273$ sec for Triton fused and $1,111$ sec for Triton fused (mem save), indicating that Triton fused (mem save) was approximately $10\%$ faster than Triton fused. Thus, the strategy of storing state vectors in reduced precision accommodates an approximately $1.4$ times larger mini-batch size, thereby enabling more effective utilization of GPU parallel computation resources. This resulted in a throughput improvement, outweighing the overhead of type conversion.
Based on these results, for a $20$-qubit HEA model including $60,000$ parameters, we estimate the training time to be approximately $15$ to $20$ hours per epoch for a data set containing tens of thousands of training samples, such as MNIST or CIFAR-10. Moreover, since our implementation is compatible with PyTorch’s standard ecosystem, it is easy to use multi-GPU distributed training (e.g., via DistributedDataParallel in PyTorch), enabling even faster training. Therefore, our proposed method enables training deeper and larger-scale QML models within a realistic timeframe.
(a) Peak memory usage
(b) Execution time
Figure 9: (a) Peak memory usage and (b) Execution time, when varying $b$ for a $20$-qubit, $1000$-layer ($d=1,000$) HEA with a batch size of one
Table 1: Performance comparison for minimizing the sum of expectation values at the final state of an HEA with 2020-qubit, $1,000$ layers ($60,000$ parameters) for $1,000$ samples
Implementation Mini-batch size Block size Total execution time (sec) Peak memory (GB)
Triton Fused 60 10 1,273 94.456
Triton Fused(mem save) 85 20 1,111 95.378
5 Conclusion
In this paper, we proposed fusing quantum gates in the forward and backward paths to improve the throughput for the classical simulation of QML and VQAs, whose performance is often limited by memory bandwidth. In our proposed method, instead of preserving all intermediate states, we store only the output state vectors of fused variational gates, omitting the storage of intermediate states within each fused gate and recomputing them in registers. This approach allows us to reduce peak memory usage and communication with global memory, achieving a significant improvement in throughput. Specifically, we achieved an approximately $20$ times higher throughput by our proposed method, compared to the PyTorch native implementation for a typical HEA. In particular, on a midrange consumer GPU, the RTX 5070, we achieved an approximately $30$ times higher throughput. This result indicates that our proposed method is particularly effective on hardware with limited memory bandwidth. These results demonstrate that our approach significantly lowers the hardware barrier, enabling highly efficient classical simulations of QML and VQAs even on a single consumer GPU.
Furthermore, by combining our proposed method with gradient checkpointing, we verified the practical training of a deep QML model. First, to optimize the checkpoint block size (i.e., the number of HEA layers to be included in the checkpoint block), we measured the mean execution time and memory usage with a batch size of one. The results showed that the execution time is almost constant regardless of the checkpoint block size, whereas the memory usage is highly consistent with our theoretical analysis. In addition, we found that memory usage can be reduced by $30\%$ by temporarily storing the state vectors as bfloat16 converted from float32 (complex64). Based on these results, we estimate that we could train a $20$-qubit QML model with $60,000$ parameters on tens of thousands of training samples, such as MNIST or CIFAR-10, within $15$ to $20$ hours per epoch. Also, since our implementation is compatible with standard PyTorch functions, we can further accelerate training by distributing samples over multiple GPUs using utilities in PyTorch, such as DistributedDataParallel. Therefore, our combined approach enables the training of deep QML models on large datasets within a realistic timeframe, which would previously require a supercomputer or a multi-GPU cluster.
Although we focused on accelerating the training of QML models, our proposed method is expected to accelerate VQA training as well, since the performance of training VQAs in classical simulation is often limited by memory bandwidth. In addition, our implementation supports double-precision (float64) calculations, making it a powerful tool for theoretical research requiring higher numerical precision, such as quantum chemistry and the analysis of the barren plateau.
We expect that our proposed method will significantly contribute to advancing the study of QML and VQAs by accelerating the training of wider and deeper quantum circuits with larger samples, facilitating the exploration of useful quantum applications and the verification of learning theories.
Funding
This work was supported by JSPS KAKENHI Grant Number JP25K21267.
Code Availability
The code used in this study will be available at GitHub (https://github.com/puyokw/triton_qml).
Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
During the preparation of this work, we used Gemini in order to debug and refactor our code and to improve our English expressions.
Appendix A Analysis of the acceleration effect of classically simulating a single-layer HEA
To facilitate the performance comparison between our proposed method and the PyTorch native implementation presented in Fig. 8, Fig. 10 illustrates the relative throughput improvement of our proposed method compared to the PyTorch native implementation.
From Fig. 10, we observe that the mid-range GPU RTX 5070 benefits most from our proposed method because of its relatively lower memory bandwidth, compared to the high-end GPU RTX 4090 and the data center GPU GH200. This result suggests that while the performance of the PyTorch native implementation is strongly limited by memory bandwidth, our proposed method effectively mitigates this bottleneck by reducing the number of accesses to global memory via gate fusion. We also observe that the Triton fused (mem save) implementation exhibits a slightly lower acceleration ratio than the Triton fused implementation due to the overhead of type conversion. Nevertheless, it remains significantly faster than the PyTorch native implementation.
Furthermore, the periodic fluctuations are observed in Fig. 10 (with local peaks at $n=12,15,18,21,24$), which arise from the grouping strategy of gate fusion. Since we fuse single-qubit gates acting on groups of three adjacent qubits, the computational efficiency is maximized when the number of qubits is a multiple of three.
Figure 10: Throughput improvement ratio of our proposed method relative to PyTorch native implementation, based on the results in Fig. 8, for a single-layer ($d=1$) of HEA in Fig. 7.
Appendix B Details of other fused gate implementations
Here, we describe the implementation of fused gates, which were not detailed in the main text.
B.1 The implementation of fused CZ gates and fused CNOT gates
Since a CNOT gate or a CZ gate is a unitary matrix ($U=U^{\dagger}$) with no variational parameters, we can update the quantum states and adjoint states using the same kernel function for both the forward and backward paths. Here, we describe the implementation for the forward path.
Let $c$ be the index of the control qubit and $t$ be the index of the target qubit. As we explained in Sec. 3.1, applying a CZ gate acting on the $c$-th control qubit and the $t$-th target qubit corresponds to multiplying the probability amplitude $a_{k}$ by 1-1 only when both the $c$-th and $t$-th bits of the index $k$ are $1$. Similarly, for the implementation of fused multiple CZ gates, let $c_{j}$ and $t_{j}$ denote the control and target qubits for the $j$-th gate in the fused operation. Each thread first counts the number of CZ gates where both the $c_{j}$-th and $t_{j}$-th bits of the index $k$ are $1$. If this total count is odd, the thread multiplies the corresponding probability amplitude $a_{k}$ by 1-1.
Specifically, we implemented this using Triton [triton2019], as shown in Listing 1. First, we prepare a bitmask (cz_mask) where only the $c_{j}$-th and $t_{j}$-th bits are set to $1$. For each thread index (offset), we check if the bitwise-AND between the index and the mask (cz_mask & offset) equals the mask itself (cz_mask). By accumulating these boolean evaluation results across all individual CZ gates within the fused operation via an XOR operation to compute the parity, we can determine whether to flip the sign of the amplitude. By applying this calculation to all $2^{n}$ amplitudes, we can update the state vector for the fused CZ gates.
Listing 1: The Triton code for fused CZ gates
1 parity = tl.zeros_like(offset)
2 for i in tl.static_range(N_CZ_GATES):
3 cz_mask = tl.load(cz_masks_ptr + i)
4 parity = parity^((cz_mask & offset)==cz_mask)
5 should_flip = parity==1
6
7 flipped_real = -val_real
8 flipped_imag = -val_imag
9 final_real = tl.where(should_flip, flipped_real, val_real)
10 final_imag = tl.where(should_flip, flipped_imag, val_imag)
Although we used fused CZ gates in this paper, the efficient implementation of fused CNOT gates is also possible. Here, we describe how to implement a fused CNOT gate with out-of-place operations, since, as we mentioned before, PyTorch’s automatic differentiation requires inputs to be invariant. Also, recall that applying CNOT gate corresponds to swapping the pair of amplitudes between $a_{\ldots 1_{c}\ldots 0_{t}\ldots}$ and $a_{\ldots 1_{c}\ldots 1_{t}\ldots}$ (or between $a_{\ldots 0_{t}\ldots 1_{c}\ldots}$ and $a_{\ldots 1_{t}\ldots 1_{c}\ldots}$), so we can implement it by tracking the index. We start from the original index assigned to each thread, and trace the index to obtain the destination index. Specifically, as described in Listing 2, for the $j$-th CNOT gate in the fused operation, if the $c_{j}$-th bit of the current index is $1$, we update the current index (out_idx) by applying a bitwise XOR operation between the current index and a mask (target_mask) that has only the $t_{j}$-th bit set to $1$. By repeating this procedure across all individual CNOT gates within the fused operation, we obtain the final destination index. By allocating a new vector and writing the original probability amplitude to the resulting destination index (out_idx), we can update the quantum states as an out-of-place operation.
Listing 2: The Triton code for fused CNOT gates
1 for i in range(N_GATES):
2 idx = i.to(tl.int64)
3 control_mask = tl.load(control_masks_ptr + idx)
4 target_mask = tl.load(target_masks_ptr + idx)
5 condition = (out_idx & control_mask) == control_mask
6 out_idx = tl.where(condition, out_idx^target_mask, out_idx)
B.2 Efficient implementation of calculating expectation values
Here, we describe our efficient implementation for calculating the expectation values, especially for the forward path, because key implementation for the backward path, given by Eq. (4), is largely shared with the forward path.
To efficiently calculate the expectation values, we use a matrix-free approach without explicitly storing intermediate state vectors or operator matrices. As shown in Listing 3, first, to prepare the bra vector elements, each thread loads the probability amplitude of the final state $|\psi_{M}\rangle$ corresponding to its assigned index within the batch (in_batch_idx) and computes its complex conjugate.
Next, we calculate the ket vector elements of $O|\psi_{M}\rangle$ using the relation of $Y=iXZ$. Since the Pauli operators $X$ and $Y$ cause bit flips, we determine the indexes of the probability amplitudes to be loaded using a bitwise XOR: target_idx = in_batch_idx ˆ x_mask, where x_mask is a bitmask representing the target qubits for the X and Y operators in $O$. We then load the probability amplitude $a_{\text{target\_idx}}$, and apply the phase shifts by $Y$ and $Z$. In our implementation, we prepare a bitmask (z_mask) representing the target qubit indexes for $Y$ and $Z$, and calculate the bitwise AND between z_mask and target_idx. If the parity of this result (target_idx&z_mask) is odd, we multiply the amplitudes by 1-1. Subsequently, we multiply the amplitudes by $i=\sqrt{-1}$ to the power of Y_COUNT (the total number of $Y$ in $O$). Note that we provide this count Y_COUNT (specifically its modulo $4$, Y_COUNT_MOD4) as a compile-time constant to the Triton compiler in order to resolve the conditional branches during compilation, thereby avoiding the runtime overhead associated with branch divergence.
Finally, we compute the product of the obtained elements from the bra and ket vectors in each thread. After computing a reduction sum within each block, we calculate the global sum using atomic_add operation. In this way, we avoid large-scale matrix multiplications and achieve efficient calculation of the expectation values with minimized memory accesses.
Listing 3: The Triton code for calculating expectation values
1 in_batch_idx = offset.to(tl.int64)
2 psi_ptr = state_vec_ptr + 2*in_batch_idx + 2 * transforms_per_batch * batch_idx
3 psi_real = tl.load(psi_ptr, mask=mask, other=0.0)
4 psi_imag = tl.load(psi_ptr + 1, mask=mask, other=0.0)
5 bra_real = psi_real
6 bra_imag = -psi_imag
7
8 target_idx = in_batch_idx ^ x_mask
9 ket_ptr = state_vec_ptr + 2*target_idx + 2 * transforms_per_batch * batch_idx
10 ket_real = tl.load(ket_ptr, mask=mask, other=0.0)
11 ket_imag = tl.load(ket_ptr + 1, mask=mask, other=0.0)
12
13 # applying multi Z gates
14 parity = _parity_jit(target_idx & z_mask)
15 is_odd = parity == 1
16 ket_real = tl.where(is_odd, -ket_real, ket_real)
17 ket_imag = tl.where(is_odd, -ket_imag, ket_imag)
18 # (multi X gates have already applied)
19 if Y_COUNT_MOD4 == 0: # * 1
20 pass
21 elif Y_COUNT_MOD4 == 1: # * i
22 # (a+bi)*i = -b+ai
23 ket_real, ket_imag = -ket_imag, ket_real
24 elif Y_COUNT_MOD4 == 2: # * -1
25 ket_real = -ket_real
26 ket_imag = -ket_imag
27 elif Y_COUNT_MOD4 == 3: # * -i
28 # (a+bi)*(-i) = b-ai
29 ket_real, ket_imag = ket_imag, -ket_real
30
31 elem_exp_val, _ = _multiply_complex_numbers(bra_real, bra_imag, ket_real, ket_imag) # real, imag
32 block_exp_val = tl.sum(elem_exp_val, axis=0)
33 tl.atomic_add(output_exp_val + batch_idx, block_exp_val, mask=(tl.arange(0, BLOCK_SIZE) == 0))
Appendix C Memory behavior analysis of the PyTorch native implementation using $100$-layer HEA
In this section, to complement the experiments in Sec. 4.3, we describe the results of a classical simulation for $100$ HEA layers to verify the memory behavior of the PyTorch native implementation, which could not be evaluated in the large-scale numerical experiments in Sec. 4.3 due to excessive memory requirements. First, we verify our analytical model described in Sec. 3.4, especially for the PyTorch native implementation, using a smaller model consisting of $100$ HEA layers with $20$ qubits and $6,000$ parameters. In this numerical experiment, we also performed the task of minimizing the sum of expectation values for the observable $O=IXYZ\ldots IXYZ$ (with $IXYZ$ repeated $5$ times) at the final states, made by applying the quantum circuit to random initial quantum states for a $20$-qubit system.
Similar to Sec. 4.3, to determine the optimal size of $b$, we first measured the mean execution time and peak memory usage as we varied the checkpoint block size $b$ with a fixed batch size of $1$.
Analyzing the behavior of peak memory usage: As shown in Fig. 11(a), we observe that the peak memory usage for the PyTorch native implementation increases monotonically with $b$. This behavior is consistent with the theoretical peak memory usage. By substituting $l_{\text{var}}=60,l_{\text{const}}=20,d=100,M_{\text{sv}}=8\times 2^{20}$ (bytes) into Eq. (5), the peak memory usage is estimated as
$M_{\text{total}}\simeq\left(80b+\frac{100}{b}\right)\times 8\times 2^{20},$
since PyTorch’s automatic differentiation stores the input quantum states for all gates, and each HEA layer consists of $60$ single-qubit gates and $20$ CZ gates ($80$ gates in total). Fig. 11(a) shows that the theoretical values agree well with the measured ones. The peak memory usage is minimized at $b^{*}=\sqrt{\frac{100}{80}}\simeq 1.1$. This analysis aligns well with the observation that peak memory usage is minimized when $b=1$.
Similarly, the theoretical peak memory usage for the Triton fused implementation is represented by
$M_{\text{total}}\simeq(7b+\frac{100}{b})\times 8\times 2^{20},$
since we obtain $\lceil\frac{l_{\text{var}}}{\alpha}\rceil=7$ by substituting $l_{\text{var}}=60,\alpha=9,d=100,M_{\text{sv}}=8\times 2^{20}$ into Eq. (7). From this theoretical peak memory usage, the optimal block size is $b^{*}=3.7$, and the measured minimum peak memory usage is $528$ MB at $b=4$. Also, the theoretical peak memory usage for Triton fused (mem save) is
$M_{\text{total}}\simeq(3.5b+\frac{100}{b})\times 8\times 2^{20}.$
The optimal block size for Triton fused (mem save) is $b^{*}=5.3$, and the measured minimum peak memory usage is $406$ MB at $b=5$. Thus, we can see that the theoretical values for peak memory usage are highly consistent with the measured values for the Triton fused and Triton fused (mem save) implementations. Note that a slight discrepancy between theoretical and measured values is observed when memory usage is below 1 GB, primarily due to the memory overheads such as CUDA contexts.
From Fig. 11(b), we also see that the computation time remains almost constant regardless of the checkpoint block size $b$.
Comparing the training time for $1$ epoch: Here, we compare the training times among the three implementations. For the checkpoint setting, based on the results of the above numerical experiments, we set the checkpoint block size $b$ to $b=1$ (every HEA layer) for the PyTorch native implementation, $b=4$ for Triton fused, and $b=5$ HEA layers for Triton fused (mem save). We performed the task of minimizing the sum of expectation values for the observable $O=(IXYZ)^{\otimes 5}$ at the final state by applying $100$ HEA layers to $1,000$ randomly initialized quantum states with a $20$-qubit system over $1$ epoch.
As shown in Table 2, we set the mini-batch size to $60$ for the PyTorch native implementation, $180$ for Triton fused, and $250$ for Triton fused (mem save) so as to use approximately $90$ GB of GPU memory for each. The total execution times for Triton fused and Triton fused (mem save) are almost the same, and both are $18$ times faster than the PyTorch native implementation. This drastic acceleration is largely contributed to by accelerating the computation itself through gate fusion and increasing the mini-batch size by three to four times through our memory optimization technique. Therefore, this numerical experiment demonstrates that our theoretical memory analysis is valid, especially for the PyTorch native implementation, and that our proposed method is effective for accelerating the training of QML and VQAs in medium-scale quantum circuits.
(a) Peak memory usage
(b) Execution time
Figure 11: (a) Peak memory usage and (b) Execution time, when running a $100$-layer HEA with $20$ qubits and a batch size of $1$, varying the checkpoint block size $b$
Table 2: Results of training the one epoch for minimizing the sum of expectation values at the final state, made by applying a $20$-qubit, $100$-layer ($6,000$ parameters) HEA to $1,000$ randomly initialized states
Implementation Mini-batch size Block size Total execution time (sec) Peak memory (GB)
PyTorch Native $60$ 11 $1,300$ $89.658$
Triton Fused $180$ 44 $67.81$ $89.993$
Triton Fused(mem save) $250$ 55 $67.14$ $91.293$
[…]